OpenStudy (anonymous):

Russel Smith Pty is a small private start-up company that initially is involved in buying and selling of goods. During the first few days of the company's existence the following transactions occur. The owner opens a business bank account and deposits $20,000 into it on the 3rd May in order to commence business. $6,000 of the deposit is from the personal funds of the owner and $14,000 is a loan from a family member who expects to be fully repaid with interest within two years. On 4th May, the company purchases a vehicle for $5,000 paying by cheque. On 5th May, the company purchased invent

OpenStudy (anonymous):

purchased inventory (stock to sell) for $3,000 on 30 days credit. On 6th May, $2,000 of the loan is repaid. On 10th May, the owner adds $4,000 of his funds to the business bank account. On 11th May, the business manages to sell all most all of its inventory for $4,500. This is paid for partly by a cheque for $2,500 and the remainder on credit of 30 days. $500 worth of stock could not be sold at this time. I need to come up with balance sheets for the 3rd May, 5th May and 11th of May, and an income statement from the 3rd to the 11th of May.

OpenStudy (anonymous):

My attempt: Income statement is pretty simple: Income: 4,500 Expenditure: 3,000 So Net Profit = 4,500 - 3,000 = 1,500 BALANCE SHEET MAY 3rd: Assets: 20,000 Liabilities: 14,000 Funds: 6,000 BALANCE SHEET MAY 5th: Assets: 23,000 Liabilities: 14,000 + 3,000 = 17,000 Funds: 6,000 BALANCE SHEET MAY 11th: Assets: 23,000 + 500 + (4,500 - 2,500) + 4,000 = 26,500 Liabilities: 12,000 + 3,000 = 15,000 Funds: 6,000 + 4,000 + Net Profit = 6,000 + 4,000 + 1,500 = 11,500

OpenStudy (anonymous):

am I correct?